OpenStudy (anonymous):

81 27 9 3 1 ? a)0 b)3 c) 1/3 d)1/9

OpenStudy (anonymous):

what to do witht his..?

OpenStudy (anonymous):

complete the series.. whats the next num after 1?

OpenStudy (anonymous):

its dividing by 3

OpenStudy (anonymous):

1st term is 81 2nd term is 27 i.e. 81/3=27 3rd term is 9 i.e. 27/3=9 4th term is 3 i.e. 9/3=3 5th term is 1 i.e. 3/3=1 now tell mw what will be 6th on same steps..??

OpenStudy (anonymous):

Just divide 1 by 3 and you get your answer

OpenStudy (anonymous):

1/3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes..its 1/3

OpenStudy (anonymous):

ohh.. thanx.. but can u tell me how can we figure out the difrnc of adding or multiplying or divinding?

OpenStudy (anonymous):

for multiplying divide 2nd term by 1st for addtion subtract 2nd term by 1st term..

OpenStudy (anonymous):

i didnt get it..

OpenStudy (anonymous):

From Mathematica: \[\text{Table}\left[3^{5-n},\{n,0,8\}\right]\to \left\{243,81,27,9,3,1,\frac{1}{3},\frac{1}{9},\frac{1}{27}\right\} \]

OpenStudy (anonymous):

c) 1/3